OpenStudy (anonymous):

Which is a solution to the equation 4x + y = –6? A. (–2, –2) B. (2, –2) C. (–2, 2) D. (2, –4)

OpenStudy (anonymous):

What do you think is the solution?

OpenStudy (anonymous):

To find the solution, just put each point into the equation and if equals -6 then you know it is a solution. I will do the first one. A (-2, -2) 4x + y = –6 4(-2) + (-2) = –6 -8 + -2 = -6 -10 = - 6 -10 does not = -6 so the point A is not a solution. Do the next one, which is B.

OpenStudy (anonymous):

Lets do B (2, -2) B (2, -2) 4x + y = –6 4(2) + (-2) = –6 8 + -2 = -6 6 = - 6 6 does not = -6 so B is not a solution

OpenStudy (anonymous):

Argh my reply deleted itself right after I just finished it plus I was trying to figure out how to do it sorry I took so long. My answer seemed wrong anyway I came up with 14 as an answer...but thank you for guiding me through B.

OpenStudy (anonymous):

It's C because 4(-2)+2=-6 -8+2=-6 -6=-6

OpenStudy (anonymous):

Correct!!!!!! Good Job

OpenStudy (anonymous):

Thanks I couldn't have figured it out without you! Thank you so much for your help and time. ^^